# Prompting Large Language Models for Recommender Systems: A Comprehensive Framework and Empirical Analysis

### v1

@@ 4.3.1. LLMs Compared to Traditional Recommenders (RQ1)
(b) Amazon-Books
Figure 5. The recommendation performance of LLMs w.r.t. the context length.
$\bullet$ Parameter scale. It is widely recognized that the larger the parameter size, the more powerful the LLMs (Zhao et al., 2023; Kang et al., 2023; Hou et al., 2023), the same applies in the field of recommender systems. To compare the effect and efficiency of LLMs w.r.t. the parameter scale, we compare the recommendation performance and inference time of LLaMA (Touvron et al., 2023a), Vicuna (Chiang et al., 2023), LLaMA2 (Touvron et al., 2023b), and LLaMA2-chat at different parameter scales in Fig. 4. As the scale of parameters enlarges, the recommendation performance and inference time of LLMs steadily increases, and both datasets (Fig. 4(a) and Fig. 4(b)) have consistent conclusions. Therefore, it is necessary to consider the trade-off between performance and efficiency when choosing the parameter scale. Moreover, the performance improvement on the increasing scale of LLaMA2 is more significant than that of LLaMA, indicating that the scale effect of LLMs has to do with the capabilities of the base model.
$\bullet$ Context length. Different LLMs have different maximum input limitations (Zhao et al., 2023; Touvron et al., 2023a, b), and a longer context input means LLMs can accommodate more historical items for recommendation. However, it remains to be explored whether the maximum input length of LLMs will affect the recommendation results when the length limitation is not exceeded. Therefore, we conduct experiments to investigate the differences in the recommendation performance between the two length versions of ChatGPT (4K and 16K) and ChatGLM3 (2K and 32K). As shown in Fig. 5, expanding the length limitation of LLMs does not necessarily mean the better recommendation performance, while there is a slight decrease in NDCG@10. Furthermore, when the maximum input of LLMs remains unchanged, increasing the historical input of users results in insignificant gains as shown in Fig. 3. Therefore, the key to the recommendation problem is to enable LLMs to effectively utilize the information within the limited context (Lin et al., 2023b; Yao et al., 2023), and a suitable context length selection for LLMs as recommender systems is worthy of deep consideration.
Observations on Recommendations w.r.t. Four Aspects of LLMs As for the public availability, closed-source models outperform the open-source models in terms of the recommendation performance, but have poorer flexibility. As for the model architecture, different frameworks are adapted to different recommendation tasks and fine-tuning strategies, while LLMs with the casual decoder architecture are still mainstream. As for the parameter scale, the larger the parameter scale, the better the recommendation ability. As for the context length, a longer maximum context length leads to worse recommendation results.
@@
dataset ml-1m Amazon-Books
@@
4.3.3. Comparisons of Tuning Strategies for LLMs (RQ3)
Due to the fact that LLMs are not customized to recommender systems during the training process, it is insufficient to only consider the zero-shot recommendation performance in cold-start scenarios (Haleem et al., 2022; Kang et al., 2023; Liu et al., 2023d; Luo et al., 2023b). In order to explore the impact of different training strategies of LLMs on recommendations, we compare the click-through rate prediction performance of four LLaMA-based LLMs on two datasets. Specifically, we consider three training settings of LLMs, i.e., the zero-shot setting without fine-tuning, Parameter-Efficient Fine-Tuning (PEFT) setting (we use the LoRA (Hu et al., 2021) here) with a few parameters tuned, and Fine-Tuning (FT) setting with all parameters tuned, and summarize empirical conclusions from three aspects: overall performance of different settings, the impact of instruction tuning, and the impact of training data.
$\bullet$ Overall performance of different settings. As shown in Table 5, we can see that the results of fine-tuning LLMs (PEFT and FT) on only 256 samples are significantly better than the zero-shot performance in cold-start scenarios, and empirical findings are consistent across four LLMs (LLaMA, Alpaca-LoRA, LLaMA2, LLaMA2-chat) on both datasets. Furthermore, considering the two kinds of fine-tuning strategies, the performance of the fine-tuning setting is even better than that of the PEFT setting since more parameters are tuned (Hu et al., 2021). In addition to recommendation effects, training efficiency is also a performance that deserves attention. Therefore, we compare the training time of the two fine-tuning strategies on two datasets. As shown in Fig. 6, the time for parameter-efficient fine-tuning is significantly less than that for fine-tuning all parameters, which is in agreement with the existing literature (Hu et al., 2021; Zhao et al., 2023). Therefore, the balance between efficiency and effectiveness needs to be further considered.
@@ 4.3.3. Comparisons of Tuning Strategies for LLMs (RQ3)
(b) Amazon-Books
Figure 6. The comparison of training time w.r.t. the parameter-efficient fine-tuning (PEFT) and fine-tuning (FT) strategies.
@@
Figure 7. The fine-tuned recommendation performance of LLMs w.r.t. the number of sampled data, and LoRA is utilized.
@@
$\bullet$ The impact of instruction tuning. Before fine-tuning LLMs using the recommendation data, we are also concerned about whether the instruction tuning of LLMs using general data can further enhance the subsequent fine-tuning on recommendations. Therefore, we select two pairs of LLMs as controls, namely LLaMA and Alpaca-LoRA, as well as LLaMA2 and LLaMA2-chat. As for the Alpaca-LoRA model, it is the tuned version of LLaMA (Touvron et al., 2023a) using the the LoRA strategy (Hu et al., 2021) on the Alpaca dataset (Taori et al., 2023). As for the LLaMA2-chat model, it is also the updated version of LLaMA2 using the RLHF strategy (Touvron et al., 2023b) on the conversational dataset. As illustrated in Table 5 and Fig. 7, the recommendation results of Alpaca-LoRA are better than that of LLaMA, and the performance of LLaMA2-chat outforms that of LLaMA2, indicating the effectiveness of performing fine-tuning on general instructions. That is to say, enhancing the general knowledge of language models can also improve the domain capabilities of LLMs on specified recommendation tasks (Bao et al., 2023b).
@@
$\bullet$ The impact of training data. For few-shot fine-tuning strategies of LLMs, we conduct experiments to explore the impact of training samples on the recommendation performance. As shown in Fig. 7, we visualize the fine-tuning results corresponding to different numbers of training samples with the LoRA strategy, and also consider the traditional recommender SASRec (Kang and McAuley, 2018). Despite a few sampled data, LLMs such as LLaMA-7B can quickly adapt to the task of CTR predictions, achieving noticeable improvements as the number of training samples increases. With only a small number of training samples, we can stimulate the appreciable recommendation ability of LLMs, reflecting their remarkable emergence ability. However, when the training samples increase from 0 to 256, the performance of the traditional recommendation model SASRec still remains 0.5. Experimental results show that LLMs have advantages over conventional recommenders on abilities of few-shot learning and adapting (Bao et al., 2023b). In addition, the comparison without sampled data in Fig. 7 also highlights the zero-shot cold-start ability of large language models.
@@
Observations on Tuning Strategies of LLMs For LLM-based recommendations, few-shot training results are better than the zero-shot performance, and fine-tuning all parameters is more effective but less efficient than parameter-efficient fine-tuning. The instruction tuning using general data can further enhance the fine-tuning results of LLMs. In few-shot training scenarios, LLMs are more capable of adapting to recommendation tasks.
Figure 8. Case study of limitations when utilizing LLMs as recommender systems.
@@ 4.3.4. Case study of limitations (RQ4)
Despite powerful capabilities of LLMs, there are also some limitations using LLMs as recommender systems. To illustrate the limitations of LLMs for recommendation, we choose two recommendation cases of the powerful closed-source model GPT-4 in Fig. 8 to analyze the possible failure scenarios of LLMs.
@@
$\bullet$ Position bias. On the one hand, the generation results of LLMs have randomness, leading to instability in recommendation results, and the position bias is a typical manifestation. As shown in the first case of Fig. 8, the LLM places the ground-truth movie “The Mask” at the bottom of ranking results, because “The Mask” is at the end of the candidate item sequence. When the order of candidate items is adjusted, the recommendation results will also change, indicating the position bias during recommendations. Existing literature has also found the position bias in LLM-based recommendations (Hou et al., 2023; Ma et al., 2023; Zhang et al., 2023f), and methods such as bootstrapping (Hou et al., 2023) and Bayesian probability framework (Ma et al., 2023) have been proposed to calibrate unstable results. However, the instability of LLMs is still an unresolved issue.
@@
$\bullet$ Lack of domain knowledge. On the other hand, the lack of domain knowledge in recommendations can lead to misunderstandings of LLMs. From the right example in Fig. 8, it can be seen that the LLM places the second movie in the candidate list, i.e., “Three Colors: White” at the first position in the re-ranking result, which still indicates the position bias. Furthermore, the LLM places the ground-truth item “Dick” at the bottom of the ranking list. The possible reason is that relying solely on the movie title “Dick”, LLMs cannot infer that the movie is a comedy film about two teenage girls, which has a greater similarity to the recently watched movie “Showgirls” by users. This case demonstrates that LLMs may make inappropriate decisions due to a lack of domain knowledge in the field of recommender systems.
In addition to the two cases in Fig. 8, LLMs are also limited by several factors such as the inference time (Bao et al., 2023a), context length (Lin et al., 2023b), and memory cost (Kang et al., 2023). Therefore, exploring the practical application of LLMs through methods such as knowledge distillation (Sun et al., 2024; Tian et al., 2023) has both academic and industrial values.
@@
Observations on Limitations of LLMs Despite capabilities of LLMs, there are limitations when employing LLMs as recommender systems. The generation results of LLMs have randomness, leading to instability such as the position bias. The lack of domain knowledge in recommendations can lead to misunderstandings of LLMs.
@@ 5. PROMPT ENGINEERING SPECIALIZED FOR RECOMMENDER SYSTEMS
Prompt is an important medium for interactions between humans and language models, and a well-designed prompt can better stimulate the powerful capabilities of LLMs (Liu et al., 2023f; Le Scao and Rush, 2021). Improving the performance of artificial intelligence by designing and improving prompts is known as prompt engineering. When prompting LLMs as recommender systems, although the specific prompts in different studies may not be the same, the format is largely identical with only minor differences (Hou et al., 2023; Gao et al., 2023; Dai et al., 2023a; Wang et al., 2023c). In our work, we concentrate on the general framework of prompt engineering specialized for recommender systems, and summarize four key elements of prompt formats. Firstly, suitable task description is the primary condition for constructing prompts (Geng et al., 2022; Wang et al., 2023c). Secondly, the characteristic of recommendation tasks lies in the mining and utilization of the user personalized interest, and user interest modeling is the essential aspect to reflect the true intentions and preferences of users (Xi et al., 2023; Yao et al., 2023). Thirdly, the purpose of recommender systems is to provide users with appropriate items, and the selection, matching, or generation of candidate items is a worthwhile issue to consider, i.e., candidate item construction (Lin et al., 2023b). Fourthly, prompting strategies are important for eliciting the planning and reasoning abilities of LLMs, which are the key to distinguishing LLMs from other language models (Zhao et al., 2023). As shown in Figure 2(a), we refer to prompts in related work (Hou et al., 2023; Zhao et al., 2023) as the basic prompt, and further analyze the effects of task description, user interest modeling, candidate items construction and prompting strategies in the experimental part.

### v2

@@ 4.3.1. LLMs Compared to Traditional Recommenders (RQ1)
(b) Amazon Books
Figure 5. The recommendation performance of LLMs w.r.t. the context length. We compare the different context length of LLMs, and the larger the symbol in the figure, the longer the context length.
$\bullet$ Parameter scale. It is widely recognized that the larger the parameter size, the more powerful the LLMs (Zhao et al., 2023; Kang et al., 2023; Hou et al., 2024), the same applies in the field of recommender systems. To compare the effect and efficiency of LLMs w.r.t. the parameter scale, we compare the recommendation performance and inference time of LLaMA (Touvron et al., 2023a), Vicuna (Chiang et al., 2023), LLaMA2 (Touvron et al., 2023b), LLaMA2-chat and LLaMA3 (Dubey et al., 2024) at different parameter scales in Fig. 4(b). As the scale of parameters enlarges, the recommendation performance and inference time of LLMs steadily increases, and both datasets (Fig. 4(a) and Fig. 4(b)) have consistent conclusions. Therefore, it is necessary to consider the trade-off between performance and efficiency when choosing the parameter scale. Moreover, the performance improvement on the increasing scale of LLaMA2 and LLaMA3 is more significant than that of LLaMA, indicating that the scaling law of LLMs has to do with the capabilities of the base model.
$\bullet$ Context length. Different LLMs have different maximum input limitations (Zhao et al., 2023; Touvron et al., 2023a, b), and a longer context input means LLMs can accommodate more historical items for recommendation. However, it remains to be explored whether the maximum input length of LLMs will affect the recommendation results when the length limitation is not exceeded. Therefore, we conduct experiments to investigate the differences in the recommendation performance between the two length versions of ChatGPT (4K and 16K) and ChatGLM3 (2K and 32K). As shown in Fig. 5, expanding the length limitation of LLMs does not necessarily mean the better recommendation performance, while there is a slight decrease in NDCG@10. Furthermore, when the maximum input of LLMs remains unchanged, increasing the historical input of users results in insignificant gains as shown in Fig. 3. Therefore, the key to the recommendation problem is to enable LLMs to effectively utilize the information within the limited context (Lin et al., 2024a; Yao et al., 2023), and a suitable context length selection for LLMs as recommender systems is worthy of deep consideration.
Observations on Recommendations w.r.t. Four Aspects of LLMs As for the public availability, closed-source models generally outperform the open-source models in the recommendation performance, but have poorer flexibility. As for the model architecture, different frameworks are adapted to different recommendation tasks and fine-tuning strategies, while LLMs with the casual decoder architecture are still mainstream. As for the parameter scale, the larger the parameter scale, the better the recommendation performance. As for the context length, a longer maximum context length leads to worse recommendation results.
@@
Dataset MovieLens-1M Amazon Books
@@
LLaMA3-8B 0.5724 0.6577 0.6927 0.7216 0.8610 0.8746
4.3.3. Comparisons of Tuning Strategies for LLMs (RQ3)
Due to the fact that LLMs are not customized to recommender systems, it is insufficient to only consider the zero-shot recommendation performance in cold-start scenarios (Haleem et al., 2022; Kang et al., 2023; Liu et al., 2023d; Luo et al., 2023). In order to explore the impact of different training strategies of LLMs on recommendations, we compare the CTR prediction performance of five LLaMA-based LLMs on two datasets. Specifically, we consider three training settings of LLMs, i.e., the zero-shot setting without fine-tuning, Parameter-Efficient Fine-Tuning (PEFT) setting (we use the LoRA (Hu et al., 2021) here) with a few parameters tuned, and Fine-Tuning (FT) setting with all parameters tuned, and summarize empirical conclusions from three aspects: overall performance of different settings, the impact of instruction tuning, and the impact of training data.
$\bullet$ Overall performance of different settings. As shown in Table 5, we can see that the results of fine-tuning LLMs (PEFT and FT) on only 256 samples are significantly better than the zero-shot performance in cold-start scenarios, and empirical findings are consistent across four LLMs on both datasets. Furthermore, considering the two kinds of fine-tuning strategies, the performance of the fine-tuning setting is even better than that of the PEFT setting since more parameters are tuned (Hu et al., 2021). In addition to recommendation effects, training efficiency is also a performance that deserves attention. Therefore, we compare the training time of the two fine-tuning strategies on two datasets. As shown in Fig. 6, the time for PEFT is significantly less than that for fine-tuning all parameters, which is in agreement with the existing literature (Hu et al., 2021; Zhao et al., 2023). Therefore, the balance between efficiency and effectiveness needs to be further considered.
Table 6. Impact of LoRA hyper-parameters ($r$ and $\alpha$ represents the rank and scaling factor of LoRA, respectively) on the recommendation performance of LLaMA3-8B. Here, $r$ typically ranges from 4 to 64, and $\alpha$ is usually set to twice the value of $r$.
$r$ $\alpha$ MovieLens-1M Amazon Books
4 8 0.6378 0.8612
8 16 0.6380 0.8610
16 32 0.6592 0.8672
@@ 4.3.3. Comparisons of Tuning Strategies for LLMs (RQ3)
[…]
Table 7. Impact of the dropout rate, a hyper-parameter in LoRA, on the performance of LLaMA3-8B fine-tuned as a CTR recommendation model on the MovieLens-1M and Amazon Books datasets.
Dropout MovieLens-1M Amazon Books
0 0.6375 0.8593
[…]
0.2 0.6577 0.8548
0.3 0.6388 0.8516
0.4 0.6429 0.8610
0.5 0.6473 0.8420
@@
(b) Amazon Books
Figure 6. Comparison of training time w.r.t. two fine-tuning strategies, i.e., PEFT and FT. The training time of LLMs is averaged over all training epochs, measured in seconds (s). Five LLaMA-based models are compared.
@@
Figure 7. The fine-tuned performance of LLMs w.r.t. the number of sampled data, and LoRA is utilized. For LLaMA-based LLMs, we analyze the effect of sampled data on the recommendation performance (measured by NDCG@10) of few-shot learning, and compare with the traditional recommender SASRec (Kang and McAuley, 2018).
[…]
The fine-tuned performance of LLMs w.r.t. the number of sampled data, and LoRA is utilized. For LLMs, we analyze the effect of sampled data on the recommendation performance (measured by NDCG@10) of few-shot learning, and compare with SASRec (Kang and McAuley, 2018).
@@
$\bullet$ The impact of instruction tuning. Before fine-tuning LLMs using the recommendation data, we are also concerned about whether the instruction tuning of LLMs using general data can further enhance the subsequent fine-tuning on recommendations. Therefore, we select two pairs of LLMs as controls, namely LLaMA and Alpaca-LoRA, as well as LLaMA2 and LLaMA2-chat. As for the Alpaca-LoRA model, it is the tuned version of LLaMA (Touvron et al., 2023a) using the the LoRA strategy (Hu et al., 2021) on the Alpaca dataset (Taori et al., 2023). As for the LLaMA2-chat model, it is also the updated version of LLaMA2 using the RLHF strategy (Touvron et al., 2023b) on the conversational dataset. As illustrated in Table 5 and Fig. 7(b), the recommendation results of Alpaca-LoRA are better than that of LLaMA, and the performance of LLaMA2-chat outforms that of LLaMA2, indicating the effectiveness of performing fine-tuning on general instructions. That is to say, enhancing the general knowledge of language models can also improve the domain capabilities of LLMs on specified recommendation tasks (Bao et al., 2023b).
$\bullet$ The impact of hyper-parameters of LoRA. As depicted in Table 7 and Table 7, we analyze the CTR prediction performance of LLaMA3-8B with different hyper-parameters of LoRA. We can find that increasing the rank ($r$) and scaling factor ($\alpha$) generally enhances the recommendation performance, likely due to the increased model capacity with more parameters tuned. Additionally, the impact of the dropout rate varies between datasets. For MovieLens-1M, a dropout rate of 0.2 yields the highest score of 0.6577, indicating that moderate dropout enhances generalization by preventing overfitting. While for Amazon Books, a dropout rate of 0.4 results in the best performance (0.8610), suggesting that the optimal dropout rate is context-dependent. These findings underscore the importance of meticulously tuning LoRA hyper-parameters to optimize performance of LLMs in CTR recommendation tasks.
@@
$\bullet$ The impact of training data. For fine-tuning LMs with few-shot learning, we conduct experiments to explore the impact of training samples on the recommendation performance. As shown in Fig. 7(b), we visualize the fine-tuning results corresponding to different numbers of training samples with the LoRA strategy, and also consider the traditional recommender SASRec (Kang and McAuley, 2018). Despite a few sampled data, LLMs such as LLaMA-7B can quickly adapt to the task of CTR predictions, achieving noticeable improvements as the number of training samples increases. With only a small number of training samples, we can stimulate the appreciable recommendation ability of LLMs, reflecting their remarkable emergence ability. However, when the training samples increase from 0 to 256, the performance of the traditional recommendation model SASRec still remains 0.5. Experimental results show that LLMs have advantages over conventional recommenders on abilities of few-shot learning and adapting (Bao et al., 2023b). In addition, the comparison without sampled data in Fig. 7(b) also highlights the zero-shot cold-start ability of LLMs.
@@
Observations on Tuning Strategies of LLMs For LLM-based recommendations, few-shot training results are better than the zero-shot performance, and fine-tuning all parameters is more effective but less efficient than parameter-efficient fine-tuning. The instruction tuning using general data can further enhance the fine-tuning results of LLMs on specified recommendation tasks. As for parameter-efficient fine-tuning LLMs as recommenders with LoRA, increasing the rank ($r$) and scaling factor ($\alpha$) of LoRA generally enhances the recommendation performance due to the increased model capacity with more parameters tuned. In few-shot training scenarios, LLMs are more capable of adapting to recommendation tasks compared to traditional recommendation models.
(a) Case study of the position bias when leveraging GPT-4 (OpenAI, 2023) as the zero-shot re-ranker on the MovieLens-1M dataset. LLMs tend to select items in front while the ground-truth movie appears at the end of the candidate list. GPT-4 is affected by the position bias of the candidate list.
(b) Case study of the lack of domain knowledge when leveraging GPT-4 (OpenAI, 2023) as the zero-shot re-ranker on the MovieLens-1M dataset. “Dick” is a comedy film about two teenager girls, which is relevant to the movie “Showgirls”, but GPT-4 ranks “Dick” at the end of the list.
(c) Case study of misunderstanding output formats when leveraging DeepSeek-R1 (DeepSeek-AI, 2024) as the zero-shot re-ranker on the Amazon Books dataset. The candidate list provides 20 books while DeepSeek-R1 recommends 27 books, of which 10 are from historical interactions.
(d) Case study of hallucination when leveraging ChatGLM3-6B (AI, 2024) as the zero-shot re-ranker on the Amazon Books dataset. ChatGLM3-6B ranks the purchased book first, leaves out the ground-truth book, and generates two Chinese books not within the candidate list.
[…]
@@ 4.3.4. Case Study of Limitations (RQ4)
Despite powerful capabilities of LLMs, there are also some limitations using LLMs as recommender systems. To illustrate the limitations of LLMs for recommendation, we choose four recommendation cases of LLMs as zero-shot re-ranker in Fig. 8 to analyze the possible failure scenarios of LLMs.
@@
$\bullet$ Position bias. It is widely acknowledged that the generation results of LLMs have randomness, leading to instability in recommendation results, and the position bias is a typical manifestation. As shown in Fig. 8(a), the LLM places the ground-truth movie “The Mask” at the bottom of ranking results, because “The Mask” is at the end of the candidate item sequence. When the order of candidate items is adjusted, the recommendation results will also change, indicating the position bias during recommendations. Existing literature has also found the position bias in LLM-based recommendations (Hou et al., 2024; Ma et al., 2023; Zhang et al., 2023e), and methods such as bootstrapping (Hou et al., 2024) and Bayesian probability framework (Ma et al., 2023) have been proposed to calibrate unstable results.
@@
$\bullet$ Lack of domain knowledge. The lack of domain knowledge in recommendations may also lead to misunderstandings of LLMs. As shown in Fig. 8(b), the LLM places the second movie in the candidate list, i.e., “Three Colors: White” at the first position in the re-ranking result, which still indicates the position bias. Furthermore, the LLM places the ground-truth item “Dick” at the bottom of the ranking list. The possible reason is that relying solely on the movie title “Dick”, LLMs cannot infer that the movie is a comedy film about two teenage girls, which has a greater similarity to the recently watched movie “Showgirls” by users. This case demonstrates that LLMs may make inappropriate decisions due to a lack of domain knowledge in the field of recommender systems.
$\bullet$ Misunderstanding output formats. LLMs may exhibit limitations in retaining critical information regarding output formats from prompt instructions, potentially leading to responses that deviate from the intended requirements. As illustrated in Fig 8(c), while the input prompt explicitly requires the model to re-rank only 20 candidate books, the model generates 27 outputs, including 10 books from the historical interaction sequence. Furthermore, three books from the original candidate list are omitted in the output. This phenomenon directly contributes to the observed suboptimal performance in the coverage metric, as demonstrated in Table 4, where the output of LLMs frequently fails to achieve 100% coverage of the expected results.
$\bullet$ Hallucination. The hallucination problem in LLM-based recommender systems refers to the phenomenon where LLMs generate items that is not grounded in their training data or the provided input, often producing factually incorrect outputs. As depicted in Fig. 8(d), the LLM ranks the recently purchased book at the top of recommended list, and leaves out the ground-truth book. Moreover, the LLM generates two Chinese books not within the candidate list. In LLM-RS, hallucination can manifest as fabricated details of recommendation reasons, incorrect factual statements of items, or irrelevant content that deviates from the intended recommendation task.
@@
Observations on Limitations of LLMs Despite capabilities, there are limitations when leveraging LLMs as recommenders. Results of LLMs have randomness, leading to instability such as the position bias. The lack of domain knowledge in recommendations can lead to misunderstandings. LLM-based recommender systems may suffer from the hallucination problem.
@@ 5. PROMPT ENGINEERING SPECIALIZED FOR RECOMMENDER SYSTEMS
Prompt is an important medium for interactions between humans and language models, and a well-designed prompt can better stimulate the powerful capabilities of LLMs (Liu et al., 2023e; Scao and Rush, 2021). Improving the performance of artificial intelligence by designing and improving prompts is known as prompt engineering When leveraging LLMs as recommender systems, although the specific prompts in different studies may not be the same, the format is largely identical with only minor differences (Hou et al., 2024; Gao et al., 2023; Dai et al., 2023a; Wang et al., 2024c). In our work, we concentrate on the general framework of prompt engineering specialized for recommender systems, and summarize four key elements of prompt formats. Firstly, suitable task description is the primary condition for constructing prompts (Geng et al., 2022; Wang et al., 2024c). As illustrated in Table 8, we list task-specific prompts for LLMs as recommender systems in our framework. Secondly, the characteristic of recommendation tasks lies in the mining and utilization of the user personalized interest, and user interest modeling is the essential aspect to reflect the true intentions and preferences of users (Xi et al., 2023; Yao et al., 2023). Thirdly, the purpose of recommender systems is to provide users with appropriate items, and the selection, matching, and generation of candidate items is a worthwhile issue to consider, i.e., candidate items construction (Lin et al., 2024a). Fourthly, prompting strategies are important for eliciting the planning and reasoning abilities of LLMs, which are the key to distinguishing LLMs from other language models (Zhao et al., 2023). In this section, we will discuss the key components of prompts in LLM-RS.